### 10 Apr 2024 v1

# Zero-shot Logical Query Reasoning on any Knowledge Graph

Abstract
Complex logical query answering (CLQA) in knowledge graphs (KGs) goes beyond simple KG completion and aims at answering compositional queries comprised of multiple projections and logical operations. Existing CLQA methods that learn parameters bound to certain entity or relation vocabularies can only be applied to the graph they are trained on which requires substantial training time before being deployed on a new graph. Here we present UltraQuery, an inductive reasoning model that can zero-shot answer logical queries on any KG. The core idea of UltraQuery is to derive both projections and logical operations as vocabulary-independent functions which generalize to new entities and relations in any KG. With the projection operation initialized from a pre-trained inductive KG reasoning model, UltraQuery can solve CLQA on any KG even if it is only finetuned on a single dataset. Experimenting on 23 datasets, UltraQuery in the zero-shot inference mode shows competitive or better query answering performance than best available baselines and sets a new state of the art on 14 of them. The code is available: https://github.com/DeepGraphLearning/ULTRA
Graph Machine Learning, Knowledge Graphs
1 Introduction
Complex logical query answering (CLQA) generalizes simple knowledge graph (KG) completion to more complex, compositional queries with logical operators such as intersection $(\wedge)$, union $(\lor)$, and negation $(\lnot)$. Such queries are expressed in a subset of first-order logic (FOL) where existentially quantified $(\exists)$ variables and given constants comprise relation projections (or atoms), and logical operators combine projections into a logical query (graph pattern). A typical example of a logical query (Ren et al., 2023) is presented in Figure 1: $U.\exists V:\texttt{Win}(\texttt{NobelPrize},V)\land\texttt{Citizen}(\texttt{USA},V)\land\texttt{Graduate}(V,U)$ where $\texttt{Win}()$ is a relation projection, NobelPrize is a constant, and $V$ is an existentially quantified variable.
Figure 1: The inductive logical query answering setup where training and inference graphs (and queries) have different entity and relation vocabularies. We propose a single model (UltraQuery) that zero-shot generalizes to query answering on any graph with new entity or relation vocabulary at inference time.
Due to the incompleteness of most KGs, these logical queries cannot be directly solved by graph traversal algorithms. Consequently, CLQA methods have to deal with missing edges when modeling the projection operators. The vast majority of existing CLQA methods (Ren et al., 2023) predict missing edges by learning graph-specific entity and relation embeddings making such approaches transductive and not generalizable to other KGs. A few approaches (Galkin et al., 2022b) are able to generalize query answering to new nodes at inference time but still need a fixed relation vocabulary.
In this work, we focus at the hardest inductive generalization setup where queries and underlying graphs at inference time are completely different from the training graph, i.e., both entities and relations are new. Furthermore, we aim at performing CLQA in the zero-shot setting with one single model. That is, instead of adaptively finetuning a model on each target dataset, we seek to design a unified approach that generalizes to any KG and any complex query at inference time. For example, in Figure 1, the training graph describes academic entities with relations Win, Citizen, Graduate [^1] We assume the presence of respective inverse relations $\texttt{r}^{-1}$. whereas the inference graph describes music entities with relations Band Member and Plays. The query against the inference graph $U:\texttt{BandMember}(\texttt{Dire Straits},U)\wedge\texttt{Plays}^{-1}(\texttt{Guitar},U)$ involves both new entities and relations and, to the best of our knowledge, cannot be tackled by any existing CLQA method that learns a fixed set of entities or relation embeddings from the training graph.
To this end, we leverage the recent progress in foundation models for KG reasoning (Galkin et al., 2023; Gao et al., 2023) and devise UltraQuery, the first CLQA approach that generalizes to logical queries on any arbitrary KG with any entity and relation vocabulary in the zero-shot fashion without relying on any external node or edge features. UltraQuery parameterizes the projection operator by an inductive graph neural network (GNN) and implements non-parametric logical operators with fuzzy logics (van Krieken et al., 2022). The pre-trained projection operator (Galkin et al., 2023) does not learn any graph-specific entity nor relation embeddings thanks to the generalizable meta-graph representation of relation interactions, and therefore enables zero-shot generalization to any KG.
In the absence of existing datasets for our inductive generalization setup, we curate a novel suite of 11 inductive query answering datasets where graphs and queries at inference time have new entity and relation vocabularies. Experimentally, we train a single UltraQuery model on one dataset and probe on other 22 transductive and inductive datasets. Averaged across the datasets, a single UltraQuery model outperforms by 50% (relative MRR) the best reported baselines in the literature (often tailored to specific graphs) on both EPFO queries and queries with negation.
2 Related Work
Complex Logical Query Answering. To the best of our knowledge, there is no existing approach for generalizable and inductive query answering where the method is required to deal with new entitis and relations at inference time.
Due to the necessity of learning entity and relation embeddings, the vast majority of existing methods like GQE (Hamilton et al., 2018), BetaE (Ren & Leskovec, 2020), ConE (Zhang et al., 2021b), MPQE (Daza & Cochez, 2020) (and many more from the survey by Ren et al. (2023)) are transductive-only and tailored for a specific set of entities and relations. Among them, CQD (Arakelyan et al., 2021) and QTO (Bai et al., 2023) are inference-only query answering engines that execute logical operators with non-parametric fuzzy logic operators (e.g., product logic) but still require pre-trained entity and relation embedding matrices to execute relation projections (link prediction steps). We refer the interested reader to the comprehensive survey by Ren et al. (2023) that covers query answering theory, a taxonomy of approaches, datasets, and open challenges.
A few models (Zhu et al., 2022; Galkin et al., 2022b; Gebhart et al., 2023) generalize only to new entities by modeling entities as a function of relation embeddings. Gebhart et al. (2023) apply the idea of cellular sheaves and harmonic extension to translation-based embedding models to answer conjunctive queries (without unions and negations). NodePiece-QE (Galkin et al., 2022b) trains an inductive entity encoder (based on the fixed vocabulary of relations) that is able to reconstruct entity embeddings of the new graph and then apply non-parametric engines like CQD to answer queries against new entities. The most effective inductive (entity) approach is GNN-QE (Zhu et al., 2022; Galkin et al., 2022b) that parameterizes each entity as a function of the relational structure between the query constants and the entity itself. However, all these works rely on a fixed relation vocabulary and cannot generalize to KGs with new relations at test time. In contrast, our model uses inductive relation projection and inductive logical operations that enable zero-shot generalization to any new KG with any entity and relation vocabulary without any specific training.
Inductive Knowledge Graph Completion. In CLQA, KG completion methods execute the projection operator and are mainly responsible for predicting missing links in incomplete graphs during query execution. Inductive KG completion is usually categorized (Chen et al., 2023) into two branches: (i) inductive entity (inductive $(e)$) approaches have a fixed set of relations and only generalize to new entities, for example, to different subgraphs of one larger KG with one set of relations; and (ii) inductive entity and relation (inductive $(e,r)$) approaches that do not rely on any fixed set of entities and relations and generalize to any new KG with arbitrary new sets of entities and relations.
Up until recently, the majority of existing approaches belonged to the inductive $(e)$ family (e.g., GraIL (Teru et al., 2020), NBFNet (Zhu et al., 2021), RED-GNN (Zhang & Yao, 2022), NodePiece (Galkin et al., 2022a), A*Net (Zhu et al., 2023), AdaProp (Zhang et al., 2023), inter alia) that generalizes only to new entities as their featurization strategies are based on learnable relation embeddings.
Recently, the more generalizable inductive $(e,r)$ family started getting more attention, e.g., with RMPI (Geng et al., 2023), InGram (Lee et al., 2023), Ultra (Galkin et al., 2023), and the theory of double equivariance introduced by Gao et al. (2023) followed by ISDEA and MTDEA (Zhou et al., 2023) models. In this work, we employ Ultra to obtain transferable graph representations and execute the projection operator with link prediction over any arbitrary KG without input features. Extending our model with additional input features is possible (although deriving a single fixed-width model for graphs with arbitrary input space is highly non-trivial) and we leave it for future work.
3 Preliminaries and Problem Definition
We introduce the basic concepts pertaining to logical query answering and KGs largely following the formalizations introduced in Galkin et al. (2022b), Ren et al. (2023), and Galkin et al. (2023).
Knowledge Graphs and Inductive Setup. Given a finite set of entities $\mathcal{V}$ (nodes), a finite set of relations $\mathcal{R}$ (edge types), and a set of triples (edges) $\mathcal{E}=(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$, a knowledge graph $\mathcal{G}$ is a tuple $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{E})$. The simplest transductive setup dictates that the graph at training time $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R}_{\textit{train}},\mathcal{E}_{\textit{train}})$ and the graph at inference (validation or test) time $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R}_{\textit{inf}},\mathcal{E}_{\textit{inf}})$ are the same, i.e., $\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}}$. By default, we assume that the inference graph $\mathcal{G}_{\textit{inf}}$ is an incomplete part of a larger, non observable graph $\hat{\mathcal{G}_{\textit{inf}}}$ with missing triples to be predicted at inference time. In the inductive setup, in the general case, the training and inference graphs are different, $\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}}$. In the easier inductive entity (inductive $(e)$) setup tackled by most of the KG completion literature, the relation set $\mathcal{R}$ is fixed and shared between training and inference graphs, i.e., $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R},\mathcal{E}_{\textit{train}})$ and $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R},\mathcal{E}_{\textit{inf}})$. The inference graph can be an extension of the training graph if $\mathcal{V}_{\textit{train}}\subseteq\mathcal{V}_{\textit{inf}}$ or be a separate disjoint graph (with the same set of relations) if $\mathcal{V}_{\textit{train}}\cap\mathcal{V}_{\textit{inf}}=\varnothing$. In CLQA, the former setup with the extended training graph at inference is tackled by InductiveQE approaches (Galkin et al., 2022b).
In the hardest inductive entity and relation (inductive $(e,r)$) case, both entities and relations sets are different, i.e., $\mathcal{V}_{\textit{train}}\cap\mathcal{V}_{\textit{inf}}=\varnothing$ and $\mathcal{R}_{\textit{train}}\cap\mathcal{R}_{\textit{inf}}=\varnothing$. In CLQA, there is no existing approach tackling this case and our proposed UltraQuery is the first one to do so.
First-Order Logic Queries. Applied to KGs, a first-order logic (FOL) query $q$ is a formula that consists of constants Con ($\texttt{Con}\subseteq{\mathcal{V}}$), variables Var ($\texttt{Var}\subseteq{\mathcal{V}}$, existentially quantified), relation projections $R(a,b)$ denoting a binary function over constants or variables, and logic symbols ($\exists,\land,\lor,\lnot$). The answers $A_{{\mathcal{G}}}(q)$ to the query $q$ are assignments of variables in a formula such that the instantiated query formula is a subgraph of the complete, non observable graph $\hat{{\mathcal{G}}}$. Answers are denoted as easy if they are reachable by graph traversal over the incomplete graph $\mathcal{G}$} and denoted as hard if at least one edge from the non observable, complete graph $\hat{{\mathcal{G}}}$ has to be predicted during query execution.
For example, in Figure 1, a query Which band member of Dire Straits played guitar? is expressed in the logical form as $U:\texttt{BandMember}(\texttt{Dire Straits},U)\wedge\texttt{Plays}^{-1}(\texttt{Guitar},U)$ as an intersection query. Here, $U$ is a projected target variable, Dire Straits and Guitar are constants, BandMember and Plays are relation projections where $\texttt{Plays}^{-1}$ denotes the inverse of the relation Plays. The task of CLQA  is to predict bindings (mappings between entities and variables) of the target variable, e.g., for the example query the answer set is a single entity ${\mathcal{A}}_{q}=\{(U,\texttt{Mark Knopfler})\}$ and we treat this answer as an easy answer as it is reachable by traversing the edges of the given graph. In practice, however, we measure the performance of CLQA approaches on hard answers.
Inductive Query Answering. In the transductive CLQA setup, the training and inference graphs are the same and share the same set of entities and relations, i.e., $\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}}$ meaning that inference queries operate on the same graph, the same set of constants Con and relations. This allows query answering models to learn hardcoded entity and relation embeddings at the same time losing the capabilities to generalize to new graphs at test time.
In the inductive entity $(e)$ setup considered in Galkin et al. (2022b), the inference graph extends the training graph $\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}}$ but the set of relations is still fixed. Therefore, the proposed models are still bound to a certain hardcoded set of relations and cannot generalize to any arbitrary KG.
In this work, we lift all the restrictions on the training and inference graphs’ vocabularies and consider the most general, inductive $(e,r)$ case when $\mathcal{G}_{\textit{inf}}\neq\mathcal{G}_{\textit{train}}$ and the inference graph might contain a completely different set of entities and relation types. Furthermore, missing links still have to be predicted in the inference graphs to reach hard answers.
Figure 2: (a) Example of ip query answering with UltraQuery: the inductive parametric projection operator (Section 4.1) executes relation projections on any graph and returns a scalar score for each entity; the scores are aggregated by non-parametric logical operators (Section 4.2) implemented with fuzzy logics. Intermediate scores are used for weighted initializion of relation projections on the next hop. (b) The multi-source propagation issue with a pre-trained link predictor for relation projection: pre-training on simple 1p link prediction is done in the single-source labeling mode (top) where only one query node is labeled with a non-zero vector; complex queries at later hops might have several plausible sources with non-zero initial weights from intermediate results (bottom) where a pre-trained operator fails.
Labeling Trick GNNs. Labeling tricks (as coined by Zhang et al. (2021a)) are featurization strategies in graphs for breaking symmetries in node representations which are particularly pronounced in link prediction and KG completion tasks. In the presence of such node symmetries (automorphisms), classical uni- and multi-relational GNN encoders (Kipf & Welling, 2017; Veličković et al., 2018; Vashishth et al., 2020) assign different automorphic nodes the same feature making them indistinguishable for downstream tasks. In multi-relational graphs, NBFNet (Zhu et al., 2021) and A*Net (Zhu et al., 2023) apply a labeling trick by using the indicator function $\textsc{Indicator}(h,v,r)$ that puts a query vector $\bm{r}$} on a head node $h$ and puts the zeros vector on other nodes $v$. The indicator function does not require entity embeddings and such models can naturally generalize to new entities (while the set of relation types is still fixed). Theoretically, such a labeling strategy learns conditional node representations and is provably more powerful (Huang et al., 2023) than node-level GNN encoders. In CLQA, only GNN-QE (Zhu et al., 2022) applies NBFNet as a projection operator making it the only approach generalizable to the inductive $(e)$ setup (Galkin et al., 2022b) with new nodes at inference time. This work leverages labeling trick GNNs to generalize CLQA to arbitrary KGs with any entity and relation vocabulary.
4 Method
We aim at designing a single model for CLQA on any KG in the zero-shot fashion, i.e., without training on a target graph. In the CLQA literature (Hamilton et al., 2018; Ren et al., 2020; Ren & Leskovec, 2020; Arakelyan et al., 2021; Zhu et al., 2022), it is common to break down query execution into a relation projection to traverse graph edges and predict missing links, and logical operators that model conjunction, disjunction, and union. Therefore, the main challenge boils down to designing inductive projection and logical operators suitable for any (even new) entity and relation vocabulary.
4.1 Inductive Relation Projection
The vast majority of CLQA methods are inherently transductive and implement relation projections as functions over entity and relation embeddings fixed to a certain KG vocabulary, e.g., with scoring functions from KG completion methods (Hamilton et al., 2018; Arakelyan et al., 2021; Bai et al., 2023), geometric functions (Ren et al., 2020; Zhang et al., 2021b), or pure neural methods (Amayuelas et al., 2022; Wang et al., 2023). The only method inductive to new entities (Zhu et al., 2022) learns relation embeddings and uses those as a labeling trick (Section 3) for a GNN that implements the projection operator.
As fixed relation embeddings do not transfer to new KGs with new relations, in UltraQuery, we opt for Ultra (Galkin et al., 2023), an inductive approach that builds relation representations dynamically using the invariance of relation interactions. Ultra leverages theoretical findings in multi-relational link prediction (Barcelo et al., 2022; Huang et al., 2023) and learns relation representations from a meta-graph of relation interactions [^2] The meta-graph can be efficiently obtained from any KG.. The meta-graph includes four learnable edge types or meta-relations (head-to-tail, tail-to-head, head-to-head, tail-to-tail) which are independent from KG’s relation vocabulary and therefore transfer across any graph. Practically, given a graph $\mathcal{G}$} and projection query $(h,r,?)$, Ultra employs labeling trick GNNs on two levels. First, it builds a meta-graph ${\mathcal{G}}_{r}$ of relation interactions (a graph of relations where each node is a unique edge type in $\mathcal{G}$}) and applies a labeling trick to initialize the query node $r$. Running a message passing GNN over ${\mathcal{G}}_{r}$ results in conditional relation representation which are used as initial edge type features in the second, entity-level GNN. There, a starting node $h$ is initialized with a query vector from the obtained relation representations and running another GNN over the entity graph (with a final sigmoid readout) returns a scalar score in $[0,1]$ representing a probability of each node to be a tail of a query $(h,r,?)$.
Figure 2(a) illustrates the intersection-projection query execution process where each projection step is tackled by the same inductive projection operator with initialization depending on the start anchor node or intermediate variables.
The multi-source propagation issue. While it is tempting to leverage Ultra pre-trained on multiple KG datasets for relation projection, there is a substantial distribution shift (Figure 2(b)) between KG completion and CLQA. Specifically, KG completion is a special case of relation projection where the input always contains a single node. By comparison, in multi-hop complex queries, several likely nodes might have high intermediate scores and will be labeled with non-zero vectors leading to the multiple sources propagation mode where a pre-trained operator is likely to fail. To alleviate the issue, we experimentally study two strategies: (1) short fine-tuning of the pre-trained projection operator on complex queries (used in the main UltraQuery model), or (2) use the frozen pre-trained operator and threshold intermediate scores setting all scores below $0<k<1$ to zero (denoted as UltraQuery LP). The insight is to limit the propagation to one or a few source nodes, thereby reducing the discrepancy between training and test distributions.
4.2 Inductive Logical Operations
Learnable logical operators parameterized by neural nets in many CLQA approaches (Hamilton et al., 2018; Ren et al., 2020; Zhang et al., 2021b; Amayuelas et al., 2022) fit a particular embedding space and are not transferable. Instead, we resort to differentiable but non-parametric fuzzy logics (van Krieken et al., 2022) that implement logical operators as algebraic operations (t-norms for conjunction and t-conorms for disjunction) in a bounded space $[0,1]$ and are used in several neuro-symbolic CLQA approaches (Arakelyan et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023; Bai et al., 2023; Yin et al., 2023b). UltraQuery employs fuzzy logical operators over fuzzy sets ${\bm{x}}\in[0,1]^{|{\mathcal{V}}|}$ as the relation projection operator assigns a scalar in range $[0,1]$ for each entity in a graph. The choice of a fuzzy logic is often a hyperparameter although van Krieken et al. (2022) show that the product logic is the most stable. In product logic, given two fuzzy sets ${\bm{x}},{\bm{y}}$, conjunction is element-wise multiplication ${\bm{x}}\odot{\bm{y}}$ and disjunction is ${\bm{x}}+{\bm{y}}-{\bm{x}}\odot{\bm{y}}$. Negation is often implemented as $\mathbf{1}-{\bm{x}}$ where $\mathbf{1}$ is the universe vector of all ones. For second- and later $i$-th hop projections, we obtain initial node states ${\bm{h}}_{v}$ by weighting a query vector ${\bm{r}}_{i}$ with their probability score $x_{v}$ from the fuzzy set of a previous step: ${\bm{h}}_{v}=x_{v}{\bm{r}}_{i}$.
4.3 Training
Following existing works (Ren & Leskovec, 2020; Zhu et al., 2022), UltraQuery is trained on complex queries to minimize the binary cross entropy loss.
${{\mathcal{L}}=-\frac{1}{|{\mathcal{A}}_{q}|}\sum_{a\in{\mathcal{A}}_{q}}\log p(a|q)-\frac{1}{|{\mathcal{V}}\backslash{\mathcal{A}}_{q}|}\sum_{a^{\prime}\in{\mathcal{V}}\backslash{\mathcal{A}}_{q}}\log(1-p(a^{\prime}|q))}$
where ${\mathcal{A}}_{q}$ is the answer to the query $q$ and $p(a|q)$ is the probability of entity $a$ in the final output fuzzy set. UltraQuery LP uses a frozen checkpoint from KG completion and is not trained on complex logical queries.
5 Experiments
Our experiments focus on answering the following research questions: (1) How does a single UltraQuery model perform in the zero-shot inference mode on unseen graphs and queries compared to the baselines? (2) Does UltraQuery retain the quality metrics like faithfullness and identify the easy answers reachable by traversal? (3) How does the multi-source propagation issue affect the final performance?
5.1 Setup and Datasets
Datasets. We employ 23 different CLQA datasets each with 14 standard query types and its own underlying KG with different sets of entities and relations. Following Section 3, we categorize the datasets into three groups (more statistics of the datasets and queries are provided in Appendix A):
Transductive (3 datasets) where training and inference graphs are the same $(\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}})$ and test queries cover the same set of entities and relations: FB15k237, NELL995 and FB15k all from Ren & Leskovec (2020) with at most 100 answers per query.
Inductive entity $(e)$ (9 datasets) from Galkin et al. (2022b) where inference graphs extend training graphs $(\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}})$ being up to 550% larger in the number of entities. The set of relations is fixed in each training graph and does not change at inference making the setup inductive with respect to the entities. Training queries might have more true answers in the extended inference graph.
Inductive entity and relation $(e,r)$ (11 datasets): we sampled a novel suite of WikiTopics-QA datasets due to the absence of standard benchmarks evaluating the hardest inductive setup where inference graphs have both new entities and relations $(\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}})$. The source graphs were adopted from the WikiTopics datasets (Gao et al., 2023), we follow the BetaE setting when sampling 14 query types with at most 100 answers. More details on the dataset creation procedure are in Appendix A.
Implementation and Training. UltraQuery was trained on one FB15k237 dataset with complex queries for 10,000 steps with batch size of 32 on 4 RTX 3090 GPUs for 2 hours (8 GPU-hours in total). We initialize the model weights with an available checkpoint of Ultra reported in Galkin et al. (2023). Following the standard setup in the literature, we train the model on 10 query types and evaluate on all 14 patterns. We employ product t-norm and t-conorm as non-parametric fuzzy logic operators to implement conjunction $(\wedge)$ and disjunction $(\lor)$, respectively, and use a simple 1x1-x negation. For the ablation study, UltraQuery LP uses the same frozen checkpoint (pre-trained on simple 1p link prediction) with scores thresholding to alleviate the multi-source propagation issue (Section 4.1). More details on all hyperparameters are available in Appendix B.
Figure 3: Zero-shot query answering performance (MRR, higher is better) of a single UltraQuery model trained on one FB15k237 queries dataset compared to the best available baselines on 23 datasets – trainable for each transductive and inductive $(e)$ dataset, and the heuristic baseline for newly introduced inductive $(e,r)$ datasets. EPFO is the average of 9 query types with $(\wedge,\lor)$ operators, Negation is the average of 5 query types with the negation operator $(\neg)$. UltraQuery LP is an ablated version trained only on 1p link prediction. On average, UltraQuery outperforms the best baselines. More results in Table 1 and Appendix C.
Evaluation Protocol. As we train an UltraQuery model only on one FB15k237 dataset and run zero-shot inference on other 22 graphs, the inference mode on those is inductive $(e,r)$ since their entity and relation vocabularies are all different from the training set.
As common in the literature (Ren & Leskovec, 2020; Ren et al., 2023), the answer set of each query is split into easy and hard answers. Easy answers are reachable by graph traversal and do not require inferring missing links whereas hard answers are those that involve at least one edge to be predicted at inference. In the rank-based evaluation, we only consider ranks of hard answers and filter out easy ones and report filtered Mean Reciprocal Rank (MRR) and Hits@10 as main performance metrics.
Other qualitative metrics include: (1) faithfullness (Sun et al., 2020), i.e., the ability to recover easy answers reachable by graph traversal. Here, we follow the setup in Galkin et al. (2022b) and measure the performance of training queries on larger inference graphs where the same queries might have new true answers; (2) the ROC AUC score to estimate whether a model ranks easy answers higher than hard answers – we compute ROC AUC over unfiltered scores of easy answers as positive labels and hard answers as negative. (3) Mean Absolute Percentage Error (MAPE) (Zhu et al., 2022) between the number of answers extracted from model’s predictions and the number of ground truth answers (easy and hard combined) to estimate whether CLQA models can predict the cardinality of the answer set.
Baselines. In transductive and inductive $(e)$ datasets, we compare a single UltraQuery model with the best reported models trained end-to-end on each graph (denoted as Best baseline in the experiments): QTO (Bai et al., 2023) for 3 transductive datasets (FB15k237, FB15k, and NELL995) and GNN-QE (Galkin et al., 2022b) for 9 inductive $(e)$ datasets. While a single UltraQuery model has 177k parameters, the baselines are several orders of magnitude larger with a parameters count depending on the number of entities and relations, e.g., a QTO model on FB15k237 has 30M parameters due to having 2000dd entity and relation embeddings, and GNN-QE on a reference FB 175% inductive $(e)$ dataset has 2M parameters. For a newly sampled suite of 11 inductive $(e,r)$ datasets, we compare against the edge-type heuristic baseline introduced in Galkin et al. (2022b). The heuristic selects the candidate nodes with the same incoming relation as the last hop of the query. More details on the baselines are reported in Appendix B
Table 1: Zero-shot inference results of UltraQuery and ablated UltraQuery LP on 23 datasets compared to the best reported baselines. UltraQuery was trained on one transductive FB15k237 dataset, UltraQuery LP was only pre-trained on KG completion and uses scores thresholding. The no thrs. version does not use any thresholding of intermediate scores (Section 4.1). The best baselines are trainable on each transductive and inductive $(e)$ dataset, and the non-parametric heuristic baseline on inductive $(e,r)$ datasets.
Model Inductive $(e,r)$ (11 datasets) Inductive $(e)$ (9 datasets) Transductive (3 datasets) Total Average (23 datasets) EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg EPFO avg neg avg MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 Best baseline 0.014 0.029 0.004 0.007 0.328 0.469 0.176 0.297 0.468 0.603 0.259 0.409 0.196 0.276 0.105 0.173 UltraQuery 0-shot 0.275 0.375 0.153 0.242 0.321 0.479 0.156 0.291 0.335 0.467 0.132 0.260 0.301 0.428 0.152 0.264 UltraQuery LP 0-shot 0.236 0.358 0.084 0.148 0.285 0.455 0.094 0.190 0.334 0.487 0.164 0.270 0.268 0.413 0.098 0.181 UltraQuery LP no thrs. 0.213 0.317 0.069 0.121 0.255 0.405 0.083 0.169 0.296 0.433 0.143 0.237 0.240 0.367 0.085 0.155
Figure 4: Mitigation of the multi-source message passing issue (Section 4) with UltraQuery: while UltraQuery LP (pre-trained only on 1p link prediction) does reach higher 1p query performance (center right), it underperforms on negation queries (center left). UltraQuery adapts to the multi-source message passing scheme and trades a fraction of 1p query performance for much better averaged EPFO, e.g., on the 3i query (right), and negation queries performance. More results are in Appendix C.
5.2 Main Experiment: Zero-shot Query Answering
In the main experiment, we measure the zero-shot query answering performance of UltraQuery trained on a fraction of complex queries of one FB15k237 dataset. Figure 3 and Table 1 illustrate the comparison with the best available baselines and ablated UltraQuery LP model on 23 datasets split into three categories (transductive, inductive $(e)$, and inductive $(e,r)$). For each dataset, we measure the average MRR on 9 EPFO queries with projection, intersection, and union operators, and 5 negation queries with the negation operator, respectively.
Averaged across 23 datasets, UltraQuery outperforms available baselines by relative 50% in terms of MRR and Hits@10 on both EPFO and negation queries (e.g., 0.30 vs 0.20 MRR on EPFO queries and 0.147 vs 0.105 on negation queries). The largest gains are achieved on the hardest inductive $(e,r)$ datasets where the heuristic baseline is not able to cope with the task. On inductive $(e)$ datasets, UltraQuery LP outperforms the trainable SOTA GNN-QE model on larger inductive inference graphs and performs competitively on smaller inductive versions. On transductive benchmarks, UltraQuery lags behind the SOTA QTO model which is expected and can be attributed to the sheer model size difference (177k of UltraQuery vs 30M of QTO) and the computationally expensive brute-force approach of QTO that materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU. Besides, the FB15k dataset is known for years to have leakages (Toutanova & Chen, 2015) from inverse edges that large embedding models memorize and exploit. On a similar note, the graph behind the NELL995 dataset is a collection of disconnected components which is disadvantageous for GNNs.
We note a decent performance of UltraQuery LP trained only on simple 1p link prediction and imbued with score thresholding to alleviate the multi-source message passing issue described in Section 4.1. Having a deeper look at other qualitative metrics in the following section, we reveal more sites where the issue incurs negative effects.
5.3 Analysis
Figure 5: Quality analysis on 9 inductive $(e)$ and 3 transductive datasets averaged across all 14 query types. Faithfullness, MRR (left): UltraQuery successfully finds easy answers in larger inference graphs and even outperforms trained GNN-QE baselines. Ranking of easy vs hard answers, ROC AUC (center): zero-shot inference methods slightly lag behind trainable GNN-QE due to assigning higher scores to hard answers. Cardinality Prediction, MAPE (right): UltraQuery is comparable to a much larger trainable baseline QTO. In all cases, UltraQuery LP is significantly inferior to the main model.
Here, we study four aspects of model performance: the effect of the multi-source message passing issue mentioned in Section 4.1, the ability to recover answers achievable by edge traversal (faithfullness), the ability to rank easy answers higher than hard answers, and the ability to estimate the cardinality of the answer set.
The multi-source message passing effect. The pre-trained Ultra checkpoint used in UltraQuery LP is tailored for singe-source message passing and struggles in the CLQA setup on later hops with several initialized nodes (Table 1). Training UltraQuery on complex queries alleviates this issue as shown in Figure 4, i.e., while 1p performance of UltraQuery LP is higher, the overall performance on EPFO and negative queries is lacking. In contrast, UltraQuery trades a fraction of 1p single-source performance to a much better performance on negative queries (about $2\times$ improvement) and better performance on many EPFO queries, for example, on 3i queries. Besides that, we note that the zero-shot performance of both UltraQuery models does not deteriorate from the increased size of the inference graph compared to the baseline GNN-QE.
Recovering easy answers on any graph. Faithfullness (Sun et al., 2020) is the ability of a CLQA model to return easy query answers, i.e., the answers reachable by edge traversal in the graph without predicting missing edges. While faithfullness is a common problem for many CLQA models, Figure 5 demonstrates that UltraQuery almost perfectly recovers easy answers on any graph size even in the zero-shot inference regime in contrast to the best baseline. Simple score thresholding does not help UltraQuery LP to deal with complex queries as all easy intermediate nodes have high scores above the threshold and the multi-source is more pronounced.
Ranking easy and high answers. A reasonable CLQA model is likely to score easy answers higher than hard ones that require inferring missing links on the fly (Galkin et al., 2022b). Measuring that with ROC AUC, we find (Figure 5) that UltraQuery is about 20% behind the baseline due to less pronounced decision boundaries (overlapping distributions of scores) between easy and hard answers’ scores. However, due to filtering the scores when computing ranking metrics, this fact does not have a direct negative impact on the overall prediction performance.
Estimating the answer set cardinality. Neural-symbolic models like GNN-QE and QTO have the advantage of estimating the cardinality of the answer set based on the final scores without additional supervision. As shown in Figure 5, UltraQuery is comparable to the larger and trainable QTO baseline on FB15k237 (on which the model was trained) as well as on other datasets in the zero-shot inference regime. Since cardinality estimation is based on score thresholding, UltraQuery LP is susceptible to the multi-source propagation issue with many nodes having a high score and is not able to deliver a comparable performance.
6 Discussion and Future Work
Limitations. The only parameterized component of UltraQuery is the projection operator and, therefore, limitations and improvement opportunities stem from the projection operator (Galkin et al., 2023) and its interplay with the multi-hop query answering framework. For instance, new mechanisms of tackling the multi-source propagation, better pre-training strategies, and scaling might positively impact the zero-shot CLQA performance.
Conclusion and Future Work. We presented UltraQuery, the first approach for inductive zero-shot complex logical query answering on any KG that combines a parameterized, inductive projection operator with non-parametric logical operators. Alleviating the multi-source message propagation issue is the key to adapt pre-trained projection operators into the multi-hop query answering framework. UltraQuery performs comparably to or better than strong baselines trained specifically on each graph and at the same time retains key qualitative features like faithfullness and answer cardinality estimation. Having a single query answering model working on any KG, the scope for future work is vast as highlighted by Ren et al. (2023) and includes, for example, better theoretical understanding of logical expressiveness bounds, supporting more query patterns beyond simple trees (Yin et al., 2023a, b), queries without anchor nodes (Barceló et al., 2023), hyper-relational queries (Alivanistos et al., 2022), queries with numerical literals (Demir et al., 2023), or temporal queries (Lin et al., 2023).
Impact Statement
We do not envision direct ethical or societal consequences of this work. Still, models capable of zero-shot inference on any graph might be applied to domains other than those designed by the authors. Positive impacts include saving compute resources and reducing carbon footprint of training specific models tailored for each graph.
Acknowledgments
JZ and BR are supported in part by the National Science Foundation (NSF) awards, CCF-1918483, CAREER IIS-1943364 and CNS-2212160, Amazon Research Award, AnalytiXIN, and the Wabash Heartland Innovation Network (WHIN). Computing infrastructure was supported in part by CNS-1925001 (CloudBank). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
JT and ZZ are supported by Intel-Mila partnership program, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund and a NRC Collaborative R&D Project (AI4D-CORE-06). This project was also partially funded by IVADO Fundamental Research Project grant PRF-2019-3583139727. The computation resource of this project is supported by Mila [^3] https://mila.quebec/, Calcul Québec [^4] https://www.calculquebec.ca/ and the Digital Research Alliance of Canada [^5] https://alliancecan.ca/.
Appendix A Datasets
A.1 Dataset statistics
First, Table 4, Table 4, and Table 4 provide the necessary details on the graphs behind the CLQA datasets. Then, Table 7, Table 7, and Table 7 list the query statistics. Transductive datasets are BetaE (Ren & Leskovec, 2020) datasets, inductive $(e)$ datasets are adopted from Galkin et al. (2022b) where validation and test inference graphs extend the training graph. The ratio denotes the size of the inference graph to the size of the training graph (in the number of nodes), that is, ${\mathcal{V}}_{\textit{inf}}/{\mathcal{V}}_{\textit{train}}$. In the following Section A.2 we provide more details in sampling 11 new inductive $(e,r)$ datasets WikiTopics-CLQA.
Table 2: Graph in transductive datasets (3) from Ren & Leskovec (2020). Inverse triples and edge types are included in the splits. Train, Valid, Test denote triples in the respective set.
Dataset Entities Rels Train Valid Test
FB15k 14,951 2,690 966,284 100,000 118,142
FB15k237 14,505 474 544,230 35,052 40,876
NELL995 63,361 400 228,426 28,648 28,534
Table 3: Graphs in inductive $(e)$ datasets (9) from Galkin et al. (2022b). Inverse triples and edge types are included in all graphs. Validation and Test splits contain an inference graph $({\mathcal{V}}_{\textit{inf}},{\mathcal{E}}_{\textit{inf}})$ which is a superset of the training graph with new nodes, and missing edges to predict (Valid and Test, respectively).
Ratio, % Rels Training Graph Validation Graph Test Graph
Entities Triples Entities Triples Valid Entities Triples Test
106% 466 13,091 493,425 13,801 551,336 10,219 13,802 538,896 8,023
113% 468 11,601 401,677 13,022 491,518 15,849 13,021 486,068 14,893
122% 466 10,184 298,879 12,314 413,554 20,231 12,314 430,892 23,289
134% 466 8,634 228,729 11,468 373,262 25,477 11,471 367,810 24,529
150% 462 7,232 162,683 10,783 311,462 26,235 10,782 331,352 29,755
175% 436 5,560 102,521 9,801 265,412 28,691 9,781 266,494 28,891
217% 446 4,134 52,455 9,062 227,284 30,809 9,058 212,386 28,177
300% 412 2,650 24,439 8,252 178,680 27,135 8,266 187,156 28,657
550% 312 1,084 5,265 7,247 136,558 22,981 7,275 133,524 22,503
Table 4: Graphs in the newly sampled inductive entity and relation $(e,r)$ WikiTopics-CLQA datasets (11). Triples denote the number of edges of the graph given at training, validation, or test. Valid and Test denote triples to be predicted in the validation and test sets in the respective validation and test graph.
Dataset Training Graph Validation Graph Test Graph
Entities Rels Triples Entities Rels Triples Valid Entities Rels Triples Test
Art 10000 65 27262 10000 65 27262 3026 10000 65 28023 3113
Award 10000 17 23821 10000 13 23821 2646 10000 17 25056 2783
Education 10000 19 14355 10000 19 14355 1594 10000 19 14193 1575
Health 10000 31 15539 10000 31 15539 1725 10000 31 15337 1703
Infrastructure 10000 37 21990 10000 37 21990 2443 10000 37 21646 2405
Location 10000 62 85063 10000 62 85063 9451 10000 62 80269 8917
Organization 10000 34 33325 10000 34 33325 3702 10000 34 31314 3357
People 10000 40 55698 10000 40 55698 6188 10000 40 58530 6503
Science 10000 66 12576 10000 66 12576 1397 10000 66 12516 1388
Sport 10000 34 47251 10000 34 47251 5250 10000 34 46717 5190
Taxonomy 10000 59 18921 10000 59 18921 2102 10000 59 19416 2157
Table 5: Statistics of 3 transductive datasets
Split Query Type FB15k FB15k-237 NELL995 Train 1p/2p/3p/2i/3i 273,710 149,689 107,982 2in/3in/inp/pin/pni 27,371 14,968 10,798 Valid 1p 59,078 20,094 16,910 Others 8,000 5,000 4,000 Test 1p 66,990 22,804 17,021 Others 8,000 5,000 4,000
Table 6: Statistics of 9 inductive $(e)$ datasets.
Ratio Graph 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni 106% training 135,613 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 6,582 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 1,000 1,000 1,000 1,000 1,000 test 5,446 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 1,000 1,000 1,000 1,000 1,000 113% training 115,523 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 10,256 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 1,000 1,000 1,000 1,000 1,000 test 9,782 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 1,000 1,000 1,000 1,000 1,000 121% training 91,228 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 12,696 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 5,000 5,000 5,000 5,000 5,000 test 14,458 10,000 10,000 10,000 10,000 10,000 10,000 10,000 10,000 5,000 5,000 5,000 5,000 5,000 133% training 75,326 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 15,541 50,000 50,000 50,000 50,000 50,000 50,000 20,000 20,000 5,000 5,000 5,000 5,000 5,000 test 15,270 50,000 50,000 50,000 50,000 50,000 50,000 20,000 20,000 5,000 5,000 5,000 5,000 5,000 150% training 56,114 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 16,229 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 5,000 5,000 5,000 5,000 5,000 test 17,683 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 5,000 5,000 5,000 5,000 5,000 175% training 38,851 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 40,000 50,000 50,000 50,000 validation 17,235 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 test 17,476 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 217% training 22,422 30,000 30,000 50,000 50,000 50,000 50,000 50,000 50,000 30,000 30,000 50,000 50,000 50,000 validation 18,168 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 test 16,902 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 300% training 11,699 15,000 15,000 40,000 40,000 50,000 50,000 50,000 50,000 15,000 15,000 50,000 40,000 50,000 validation 16,189 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 test 17,105 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 550% training 3,284 15,000 15,000 40,000 40,000 50,000 50,000 50,000 50,000 10,000 10,000 30,000 30,000 30,000 validation 13,616 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000 test 13,670 50,000 50,000 50,000 50,000 50,000 50,000 50,000 50,000 10,000 10,000 10,000 10,000 10,000
Table 7: WikiTopics-CLQA statistics: the number of queries generated per query pattern for each topic knowledge graph of WikiTopics (Gao et al., 2023). Numbers are the same for both the training and inference graph. We follow the same 14 query patterns introduced by Ren & Leskovec (2020).
Topics 1p 2p 3p 2i 3i pi ip 2in 3in pin pni inp 2u up Art 3113 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Award 2783 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Education 1575 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Health 1703 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Infrastructure 2405 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Location 8000 8917 4000 8000 8000 8000 8000 1000 1000 1000 1000 1000 8000 8000 Organization 3357 8000 4000 8000 8000 8000 8000 1000 1000 1000 1000 1000 8000 8000 People 6503 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Science 1388 10000 10000 10000 10000 10000 10000 1000 1000 1000 1000 1000 10000 10000 Sport 5190 8000 4000 8000 8000 8000 8000 1000 1000 1000 1000 1000 8000 8000 Taxonomy 2157 8000 8000 8000 8000 8000 8000 1000 1000 1000 1000 1000 8000 8000
A.2 WikiTopics-CLQA
The WikiTopics dataset introduced by Gao et al. (2023) was used to evaluate link prediction model’s zero-shot performance in the inductive $(e,r)$ setting, i.e., when the test-time inference graph contains both new entities and new relations unseen in training. It grouped relations into 11 different topics, or domains, such as art, education, health care, and sport. Two graphs, $\mathcal{G}_{\textit{train}}^{(T)}$ and $\mathcal{G}_{\textit{inf}}^{(T)}$ along with the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ and $\mathcal{E}^{(T)}_{\textit{test}}$, were provided for each topic $T$, which had the same set of relations but different (potentially overlapping) set of entities. The goal was to train models on the training graphs $\mathcal{G}_{\textit{train}}^{(T)}$ of some topic $T$, and test on the inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ of an unseen topic $T^{\prime}$. The model’s validation performance was evaluated on the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ when observing training graph $\mathcal{G}_{\textit{train}}^{(T)}$ as inputs, and its test performance was evaluated on $\mathcal{E}^{(T^{\prime})}_{\textit{test}}$ when observing the test inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ as inputs. Table 4 shows the statistics of the 11 topic-specific knowledge graphs in WikiTopics.
[…]
Appendix B Hyperparameters and Baselines
Both UltraQuery and UltraQuery LP are initialized from the pre-trained Ultra checkpoint published by the authors and have the same GNN architectures with parameter count (177k). UltraQuery is further trained for 10,000 steps on complex queries from the FB15k237 dataset. Hyperparameters of UltraQuery and its training details are listed in Table 8.
UltraQuery LP employs thresholding of intermediate fuzzy set scores as one of the ways to alleviate the multi-source propagation issue (Section 4.1). Generally, the threshold is set to 0.8 with a few exceptions:
  • 0.9 in FB 550, FB 300, FB 175
  • 0.97 in NELL995
Below, we discuss the best available baselines for each dataset family.
Transductive (3 datasets): QTO (Bai et al., 2023). QTO requires 2000dd ComplEx (Trouillon et al., 2016) entity and relation embeddings pre-computed for each graph, e.g., taking 30M parameters on the FB15k237 graph with 15k nodes. Further, QTO materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples for each graph. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU.
Inductive $(e)$ (9 datasets): GNN-QE (Zhu et al., 2022). The framework of GNN-QE is similar to UltraQuery but the backbone relation projection operator is implemented with NBFNet (Zhu et al., 2021) which is only inductive to entities and still learns graph-specific relation embeddings. Parameter count, therefore, depends on the number of unique relation types, e.g., 2M for the FB 175% split with 436 unique relations.
Inductive $(e,r)$ (11 datasets): for newly sampled datasets, due to the absence of fully inductive trainable baselines, we compare against a rule-based heuristic baseline similar to the baseline in Galkin et al. (2022b). The baseline looks up candidate entities that have the same incoming relation type as the current relation projection (note that the identity of the starting head node in this case is ignored). The heuristic filters entities into two classes (satisfying the incoming relation type or not), hence, in order to create a ranking, we randomly shuffle entities in each class. This baseline is non-parametric, does not require any training, and represents a sanity check of CLQA models. Still, as shown in Galkin et al. (2022b), the baseline might outperform certain inductive reasoning approaches parameterized by neural networks.
Table 8: UltraQuery hyperparameters. $\text{GNN}_{r}$ denotes a GNN over the graph of relations $\mathcal{G}_{r}$, $\text{GNN}_{e}$ is a GNN over the original entity graph $\mathcal{G}$}.
Hyperparameter UltraQuery training
$\text{GNN}_{r}$ # layers 6
hidden dim 64
message DistMult
aggeregation sum
$\text{GNN}_{e}$ # layers 6
hidden dim 64
message DistMult
aggregation sum
$g(\cdot)$ 2-layer MLP
Learning optimizer AdamW
learning rate 0.0005
training steps 10,000
adv temperature 0.2
traversal dropout 0.25
batch size 64
training queries FB15k237
Appendix C More Results
Table 9 corresponds to Figure 3 and Table 1 and provides MRR and Hits@10 results for UltraQuery, UltraQuery LP, and Best Baseline for each dataset averaged across 9 EPFO query types and 5 negation query types. Figure 6 is the full version of Figure 4 and illustrates the performance of all 3 compared models on 9 inductive $(e)$ datasets on 14 query types together with their averaged values (EPFO avg and neg avg, respectively).
Table 9: Full results (MRR, Hits@10) of UltraQuery LP and UltraQuery in the zero-shot inference regime on transductive, entity-inductive $(e)$, and fully inductive $(e,r)$ datasets compared to the best-reported baselines averaged across 9 EPFO query types (EPFO avg) and 5 negation query types (Negation avg). UltraQuery was fine-tuned only on FB15k237 queries. The numbers correspond to Table 1 and Figure 3.
UltraQuery LP UltraQuery Best Baseline EPFO avg Negation avg EPFO avg Negation avg EPFO avg Negation avg MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 transductive datasets FB15k237 0.221 0.379 0.081 0.165 0.252 0.395 0.090 0.195 0.335 0.491 0.155 0.295 FB15k 0.523 0.678 0.341 0.507 0.525 0.668 0.239 0.437 0.740 0.837 0.492 0.664 NELL995 0.256 0.405 0.070 0.139 0.228 0.339 0.067 0.147 0.329 0.483 0.129 0.268 inductive $(e)$ datasets FB 106% 0.204 0.353 0.056 0.119 0.244 0.396 0.110 0.210 0.373 0.504 0.256 0.377 FB 113% 0.226 0.377 0.069 0.148 0.258 0.407 0.123 0.228 0.352 0.494 0.214 0.339 FB 121% 0.253 0.418 0.083 0.171 0.296 0.456 0.144 0.283 0.323 0.462 0.173 0.291 FB 133% 0.275 0.442 0.086 0.170 0.308 0.464 0.135 0.263 0.353 0.514 0.197 0.341 FB 150% 0.283 0.456 0.094 0.195 0.323 0.482 0.150 0.288 0.339 0.493 0.167 0.303 FB 175% 0.309 0.488 0.107 0.215 0.346 0.510 0.175 0.322 0.351 0.507 0.188 0.336 FB 217% 0.338 0.516 0.121 0.237 0.373 0.536 0.194 0.349 0.346 0.492 0.174 0.301 FB 300% 0.328 0.513 0.109 0.220 0.365 0.528 0.184 0.340 0.291 0.426 0.125 0.224 FB 550% 0.346 0.527 0.122 0.236 0.375 0.536 0.193 0.337 0.222 0.331 0.091 0.158 inductive $(e,r)$ datasets Art 0.207 0.342 0.055 0.112 0.237 0.333 0.087 0.136 0.016 0.031 0.006 0.014 Award 0.192 0.353 0.018 0.034 0.178 0.287 0.108 0.193 0.004 0.006 0.002 0.002 Edu 0.137 0.222 0.052 0.104 0.187 0.257 0.094 0.150 0.008 0.014 0.003 0.005 Health 0.316 0.420 0.131 0.187 0.298 0.382 0.265 0.360 0.019 0.040 0.010 0.020 Infrastructure 0.362 0.516 0.095 0.195 0.370 0.455 0.175 0.244 0.001 0.018 0.003 0.005 Location 0.388 0.540 0.178 0.306 0.434 0.538 0.240 0.374 0.007 0.017 0.001 0.002 Organization 0.137 0.259 0.002 0.042 0.215 0.355 0.120 0.253 0.005 0.008 0.001 0.002 People 0.361 0.498 0.160 0.262 0.377 0.465 0.213 0.266 0.005 0.010 0.002 0.003 Science 0.207 0.342 0.045 0.092 0.236 0.356 0.093 0.172 0.041 0.085 0.010 0.020 Sport 0.148 0.259 0.049 0.092 0.202 0.335 0.145 0.229 0.015 0.030 0.005 0.008 Taxonomy 0.178 0.256 0.119 0.196 0.275 0.348 0.211 0.306 0.034 0.066 0.009 0.017
Figure 6: Full results on 9 inductive $(e)$ datasets corresponding to Figure 4: albeit UltraQuery LP outperforms the main UltraQuery on simple 1p queries, it suffers from the multi-source propagation issue on complex queries. UltraQuery trades a fraction of 1p query performance for significantly better average performance on 9 EPFO and 5 negation query types with particularly noticeable gains on intersection and 2in, 3in queries.